OpenStudy (anonymous):

If f’(5) = 3, find the (f(x)-f(5))/(√x-√5) limx->5

OpenStudy (anonymous):

\[\LARGE \lim_{x\to5}\frac{f(x)-f(5)}{\sqrt x-\sqrt5}\] I'll just rewrite it better, if this is what you meant... (I don't know to solve it ! ...) hope someone else helps you out. Good Luck !

OpenStudy (anonymous):

multipyly top and bottom by \(\sqrt{x}+\sqrt{5}\)

OpenStudy (anonymous):

thanks kreshnik again, I am too technically uncapable to use some latex editor.

OpenStudy (anonymous):

\[ \lim_{x\to5}\frac{f(x)-f(5)}{\sqrt x-\sqrt5}\frac{\sqrt x+\sqrt 5}{\sqrt x+\sqrt 5}\] \[ \lim_{x\to5}\frac{f(x)-f(5)(\sqrt{x}+\sqrt{5})}{x-5}\] \[f'(5)(\sqrt{5}+\sqrt{5})\]

OpenStudy (anonymous):

last line since \[ \lim_{x\to5}\frac{f(x)-f(5)}{ x-5}=f'(5)\] by definition

OpenStudy (anonymous):

@satellite73 I didn't know that, I learned something new. Great job .